OpenStudy (anonymous):

Algebra 2 question? What is the solution of the following system? -3x-2y=-12 9x+6y=-9

OpenStudy (anonymous):

A. (2,1) B. no solutions C. (-2,-1) D. infinitely many solutions

OpenStudy (anonymous):

do you want step by step or just the answer?

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

ok. letss call the set of formulas A and B. lets star with we'll get Y alone to start. first subtract -3x from each side. -2Y=-12+3X

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Then divide by -2 right

OpenStudy (anonymous):

this problem will either have decimals or fractions. witch do you prefer?

OpenStudy (anonymous):

decimals

OpenStudy (anonymous):

ok. correct Y=-6+.66X now we can plug this in to B in place of Y

OpenStudy (anonymous):

9x+6(-6=+.66X)=-9 (woops)

OpenStudy (anonymous):

Then distributive property correct?

OpenStudy (anonymous):

9x+6(-6+.66X)=-9 (3rd times the charm.) ok so know we need to get X alone. we need to distribute first. 9X-36+3.96X=-9

OpenStudy (anonymous):

Then add like terms?

OpenStudy (anonymous):

yes good job. 12.96Y+36=-9 then subtract 36 12.96Y=-45

OpenStudy (anonymous):

So y = 3.472

OpenStudy (anonymous):

-3.72

OpenStudy (anonymous):

Then put it back into the first problem so -3x-2(-3.72)=-12?

OpenStudy (anonymous):

yep. how do you think that should be done?

OpenStudy (anonymous):

-3x+7.44=-12

OpenStudy (anonymous):

-3x=-19.44

OpenStudy (anonymous):

yes what next?

OpenStudy (anonymous):

x=6.48

OpenStudy (anonymous):

nice work now you have both of the answers you need.

OpenStudy (anonymous):

That equals none of my choices up there though.

OpenStudy (anonymous):

oh.....

OpenStudy (anonymous):

give me a minute to look over it.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

is x first or Y in the options?

OpenStudy (anonymous):

(X,Y)

OpenStudy (anonymous):

then i think it's A. is this a qustion from a book or did your teacher make it?

OpenStudy (anonymous):

The book

OpenStudy (anonymous):

question*

OpenStudy (anonymous):

I thought it was B :/

OpenStudy (anonymous):

oh well all i can suggest is ask your teacher to look at it. but if anything the answer is A

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

that may be sense we got some random answer.

OpenStudy (anonymous):

yw.